OpenStudy (anonymous):

Line AB contains points A (−2, 3) and B (4, 5). Line AB has a slope that is zero undefined positive negative

OpenStudy (anonymous):

A (−2, 3) and B (4, 5) \[ \frac { 5-3}{4-(-2)}=\frac{2}{6}>0 \]

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

YW